OpenStudy (avanti):

how can i find sec(0) without using a calculator?

OpenStudy (anonymous):

1/cos theatre

OpenStudy (anonymous):

cos = adjacent/hypotenuse 1/cos = hypotenuse/adjacent

OpenStudy (nurali):

sec 0 =1/(cos 0) =1/1 =1

OpenStudy (avanti):

oh that makes sense, thank you so much!

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):